b. The loan officer rates applicants for credit. Ratings are normally distributed. The mean is 240 and the standard deviation is 60. If 49 applicants are randomly chosen, what is the probability that they will have a rating between 230 and 260? Round z scores to two decimal places
Answer by stanbon(48568)

The loan officer rates applicants for credit. Ratings are normally distributed. The mean is 240 and the standard deviation is 60. If 49 applicants are randomly chosen, what is the probability that they will have a rating between 230 and 260? Round z scores to two decimal places
-----
z(230) = (230-240)/[60/sqrt(49)] = -1.1667
z(260) = (260-240)/[60/sqrt(49)] = 2.3333
------------------------------
P(230 < z < 240) = P(-1.1667 < z < 2.333) = 0.8685
-------------
Cheers,
Stan H.